rvc (rvc):

A line has a slope of 7 and contains the point (5, −9). Which equations represent the line? Choose ALL answers that are correct. A.−7x + y = −44 B.7x – y = 44 C.(y − 9) = 7(x + 5) D.(y + 9) = 7(x − 5)

rvc (rvc):

@kewlkatcutie ur question is here

rvc (rvc):



rvc (rvc):

try this okay :)

OpenStudy (anonymous):

Thank you omg you're great <3333

OpenStudy (anonymous):

But wait here's only (y − 9) = 7(x + 5) or (y + 9) = 7(x − 5) , no (y - 9) = 7(x − 5) .. is here something I'm missing ??

rvc (rvc):